# Ultraviolet Completion of a Two-loop Neutrino Mass Model

### v1

@@ 1 Introduction
The origin of neutrino masses needed to explain a variety of neutrino oscillation experiments remains a mystery today and is a subject of intense experimental and theoretical scrutiny. New physics that go beyond the Standard Model (SM) is required to explain the oscillation data, but the nature of such new physics is completely unknown. A deeper understanding of its origin could have a profound impact on the evolution of the Universe and may also reveal new symmetries and conservation laws. The seesaw mechanism, which introduces gauge singlet fermions to the SM can elegantly explain the smallness of neutrino masses [1, 2, 3, 4, 5]; however, the scale of new physics is typically close to the grand unification scale, which cannot be directly tested. This comment also holds true when gauge non-singlet scalars ($SU(2)_{L}$ triplets in type-II seesaw [6, 7, 8]) or fermions ($SU(2)_{L}$ triplets in type-III seesaw [9]) are introduced to generate small neutrino masses.
@@
The Zee-Babu model [13, 14] extends the SM with two new scalars, $\eta^{+}(1,1,1)$ and $k^{++}(1,1,2)$, with $SU(3)_{C}\times SU(2)_{L}\times U(1)_{Y}$ quantum numbers as indicated. Their Yukawa couplings, $f_{ij}L_{i}L_{j}\eta^{+}+g_{ij}e^{c}_{i}e^{c}_{j}k^{--}$, with $f_{ij}=-f_{ji}$ and $g_{ij}=g_{ji}$ ($i,j$ are family indices here), as well as a cubic scalar coupling $\mu\,(\eta^{+}\eta^{+}k^{--})$ jointly break lepton number by two units, and would generate neutrino Majorana masses via two-loop diagrams involving the exchange of these scalar fields. An $SU(5)$ embedding of these fields can be realized by the addition of a $10_{H}$ and a $50_{H}$ to the scalar sector of minimal $SU(5)$ with antisymmetric $10_{H}$ and symmetric $50_{H}$ Yukawa coupling matrices, which is what our framework adopts. Note the model has no right-handed neutrino $\nu_{R}$, which goes well with $SU(5)$ unification, which also has no $\nu_{R}$. The mass relations $m_{b}^{0}/m_{\tau}^{0}=m_{s}^{0}/m_{\mu}^{0}=m_{d}^{0}/m_{e}^{0}=1$ among the quarks and leptons predicted by the minimal $SU(5)$ model (the superscripts 0 indicate GUT scale values of the masses), while a good order of magnitude estimate, fail in detail when compared to experimental values by more than 100%. These relations can be corrected by introducing a vector-like pair of fermions in the $5+\overline{5}$ representation [24]. Gauge coupling unification is realized in this framework consistently, see Fig. 3, by choosing the mass scales of certain particles to be below the GUT scale, as listed in Table 2. A key outcome of our analysis is that, in addition to the original diagrams mediated by $\eta^{+}$ and $k^{++}$ scalars, neutrinos gain mass through diagrams mediated by their color-triplet and color-sextet counterparts [25, 26, 27], see Fig. 1. Once the experimental limits from proton decay are combined with gauge coupling unification constraints and perturbativity of Yukawa couplings, we find that both the color-neutral and the colored Zee-Babu states should have masses below $\mathcal{O}(10^{3})$ TeV, which can potentially be discovered in future collider experiments, thereby testing the proposed model.
Neutrino mass generation in the framework of $SU(5)$ unification has been studied in conjunction with gauge coupling unification by various authors. Ref. [28] has studied type-II seesaw embedding in $SU(5)$, while Ref. [29] has proposed type-III seesaw embedding. Models of radiative neutrino masses [15, 16] have also been embedded in $SU(5)$. The original Zee model [10], where the neutrino masses arise as one-loop radiative corrections, has been embedded in $SU(5)$ in Ref. [30, 31, 32]. For the embedding of other loop-induced neutrino mass models in $SU(5)$, see for example, Ref. [33, 34, 35, 36, 37, 38, 39, 40, 41]. To our knowledge the model presented here is the first attempt to embed the Zee-Babu model within $SU(5)$.
@@ 2 Embedding the Zee-Babu model in $SU(5)$
Once the GUT symmetry is broken, the submultiplets of the scalar fields will receive non-degenerate masses. Although there is no mass relations among the submultiplets of $5_{H},24_{H}$, and $10_{H}$, there exist three relations within the components of $50_{H}$ [43, 35], which read
@@
These mass relations will play crucial role in gauge coupling unification, which we discuss in Sec. 5.
@@ 3.1 Charged fermion mass
As will be shown later in Sec. 5, from the gauge coupling unification constraints, the vector-like lepton doublet should have a mass near the GUT scale. Therefore, we integrate it out, which yields the light 3×3$\times$3 × charged lepton mass matrix and the heavy vector-like lepton mass as [44]
$M_{E}^{\mathrm{light}}=\left(\mathds{1}+\frac{1}{\xi_{E_{4}}^{2}}\xi_{E}\xi_{E}^{\dagger}\right)^{-1/2}v_{5}Y_{5},$ (3.7)
@@
$M_{D}^{\mathrm{light}}=v_{5}Y_{5}^{T}\left(\mathds{1}+\frac{1}{\xi_{D_{4}}^{2}}\xi_{D}\xi_{D}^{\dagger}\right)^{-1/2},$ (3.9)
@@ 3.2 A fit to the charged fermion masses
(at $\mu=M_{GUT}$)
$m_{d}/10^{-3}$ 1.14$\pm$0.11 1.14
$m_{s}/10^{-3}$ 21.58$\pm$1.14 21.58
$m_{b}$ 0.994$\pm$0.009 0.994
$m_{e}/10^{-3}$ 0.470692$\pm$0.000470 0.469544
$m_{\mu}/10^{-3}$ 99.3658$\pm$0.0993 99.2097
$m_{\tau}$ 1.68923$\pm$0.00168 1.68976
$\theta^{\rm{CKM}}_{12}$ $0.22739\pm 0.0006$ 0.2276
$\theta^{\rm{CKM}}_{23}/10^{-2}$ $4.858\pm 0.06$ 0.0486
$\theta^{\rm{CKM}}_{13}/10^{-3}$ $4.202\pm 0.13$ 0.00421
$\delta^{\rm{CKM}}$ $1.207\pm 0.054$ 1.203
Table 1: Inputs [45] and outputs for the numerical $\chi^{2}$ analysis. This benchmark fit with a total $\chi^{2}=0.15$ shows excellent agreement with data. As an example, for this fit, we additionally demand that the heavy down-type vectorike fermion mass is $m_{B}\sim 3$ TeV, which safely evades all present collider bounds.
@@ 3.3 Flavor violation
where $V={V^{u}_{L}}^{\dagger}X_{L}$ is a $3\times 4$ matrix. Since there is no up-type VLQ, $V^{u}_{L}$ is a $3\times 3$ unitary matrix that diagonalizes the up-quark mass matrix. Therefore, the Cabibbo–Kobayashi–Maskawa (CKM) matrix can be identified with the upper $3\times 3$ block of $V$, i.e., $V_{\mathrm{CKM}}=V^{\mathrm{upper}}_{3\times 3}$. During the fitting procedure, we also demand that the CKM matrix is correctly reproduced in a way that all non-unitarity constraints satisfied. For this purpose, we parametrize $V^{u}_{L}$ as $V^{u}_{L}=V^{u}_{\mathrm{CKM-like}}(\theta_{ij}^{u},\delta^{u})\mathrm{diag}(e^{i\alpha^{u}},e^{i\beta^{u}},1)$. The input parameters in the CKM as well as the fitted values are also summarized in Table 1.
@@
where we have defined $F^{d}=X_{L}^{\dagger}X_{L}$. Deviation of $F^{d}$ from the unit matrix gives rise to flavor changing neutral current (FCNC) processes, such as meson decay and meson-antimeson oscillations. The coupling responsible for Higgs-mediated FCNC take the form [50]
@@
Eqs. (3.11)-(3.13), together with Eq. (3.21), let us compute the relevant matrices as:
@@
$BR\big{(}\mathcal{B}_{q}\rightarrow\mu^{+}\mu^{-}\big{)}=\tau_{\mathcal{B}_{q}}\frac{G^{2}_{F}}{16\pi}\bigg{(}\frac{\alpha_{\text{em}}}{\pi s^{2}_{W}}\bigg{)}^{2}f^{2}_{\mathcal{B}_{q}}m_{\mathcal{B}_{q}}m^{2}_{\mu}\sqrt{1-\frac{4m^{2}_{\mu}}{m^{2}_{\mathcal{B}_{q}}}}|\eta^{2}_{Y}|\big{|}\lambda^{t}_{qb}Y_{0}(x_{t})+\Delta^{\mathcal{B}_{q}}_{\mu\mu}\big{|}^{2}\,.$ (3.41)
By demanding that the new contribution be smaller than the SM prediction at $95\%$ C.L, the corresponding constraints yields
@@
Finally, using the flavor violating couplings quoted above, one can straightforwardly compute the decay rates of the vectorike quark [50], and we find
@@ 3.4 Neutrino mass generation
Although $\hat{Y}_{A}$ remains anti-symmetric, $\widetilde{Y}_{A}$ is no longer anti-symmetric. For the latter, it is not required by the SM gauge symmetry. Furthermore,
@@ 4 Perturbativity bounds on the Yukawa couplings
Before moving to the gauge coupling unification analysis, in this section, we study perturbativity bounds on the Yukawa couplings and find the upper bounds on the Yukawa couplings involving the Zee-Babu scalars. In particular, we find the maximum values of Yukawa couplings at the low scale demanding that they remain perturbative up to the GUT scale. For this analysis, here we assume that the color-neutral as well as the colored Zee-Babu states all have masses at the TeV scale. Then, the evolution of the gauge and the Yukawa couplings are governed by the one-loop renormalization group equations (RGEs) given by
@@ 5 Gauge coupling unification and proton decay bounds
In this section, we perform a dedicated study to the gauge coupling unification. The two-loop RGEs [61, 62, 63] for the gauge couplings can be written as
@@
Figure 4: Result from extensive Markov chain Monte Carlo (MCMC) analysis. The $1\sigma$ (dark colored band) and $2\sigma$ (light colored band) highest posterior density (HPD) intervals of the masses of all multiplets from gauge coupling unification constraints. Gray horizontal dashed line represents the approximate collider bound of $m_{i}\geq 1$ TeV on the states. Gray horizontal dotted line depicts the maximum allowed mass of $m_{i}\leq 10$ TeV on the $(\eta_{1},\chi_{1})$ states. Moreover, gray horizontal dashed-dotted line at $\mu=10^{12}$ GeV scale shows the proton decay bounds by the color-triplet states $\chi_{2}$ and $\phi_{2}$.
@@
Finally, consistent gauge coupling unification points (from the above mentioned MCMC analysis) and proton decay bounds in the $g_{\mathrm{GUT}}$ vs $M_{\mathrm{GUT}}$ plane is depicted in Fig. 5. The corresponding proton decay bound is computed as follows. The decay rate for the dominant gauge boson mediated decay $p\rightarrow e^{+}\pi^{0}$ is given as [71, 72, 73]
@@
Here $f_{\pi}=130.2\,\mathrm{MeV}$ is the pion decay constant $D=0.804$ and $F=0.463$ are chiral Lagrangian parameters  [74, 75], $\alpha_{H}=-0.01257\,\mathrm{GeV}^{3}$ is the hadronic matrix element [76], $R_{L}=1.35$ is the long-distance renormalization factor of the four-fermion operators [77], $A_{SL}$ and $A_{SR}$ are the short distance factors with $A_{SL}\simeq A_{SR}\simeq 2$ [78], $M_{X}$ is the common mass of the $(X,Y)$ gauge bosons, and $\alpha_{\rm GUT}=g_{\rm GUT}^{2}/(4\pi)$ is the unified coupling at the GUT scale. Note that the value of $\alpha_{\rm GUT}\simeq 1/14$ is significantly larger than the corresponding value in typical non-supersymmetric unified theories where it is $\sim 1/40$. This increased $\alpha_{\rm GUT}$ results in the shortening of proton lifetime by about an order of magnitude. For the gauge coupling unification study, we have identified the unification scale $M_{\mathrm{GUT}}$ with the masses of the gauge bosons, i.e., $M_{\mathrm{GUT}}=M_{X,Y}$. Using Eq. (5.62), in Fig. 5, the current (future) proton decay bound of $\tau_{p}(p\to e^{+}\pi^{0})>2.4\times 10^{34}$ ($7.8\times 10^{34}$) yrs from Super-Kamiokande [79] (Hyper-Kamiokande [23]) is shown with gray shaded region (red dashed line). While, a large part of the parameter space is ruled out by the present experimental limits, intriguingly, 10 yrs operation of Hyper-Kamiokande will likely probe the entire available parameter space of the proposed model.

### v2

@@ 1 Introduction
The origin of neutrino masses needed to explain a variety of neutrino oscillation experiments remains a mystery today and is a subject of intense experimental and theoretical scrutiny. New physics that goes beyond the Standard Model (SM) is required to explain the oscillation data, but the nature of such new physics is completely unknown. A deeper understanding of its origin could have a profound impact on the evolution of the Universe and may also reveal new symmetries and conservation laws. The seesaw mechanism, which introduces gauge singlet fermions to the SM can elegantly explain the smallness of neutrino masses [1, 2, 3, 4, 5]; however, the scale of new physics is typically close to the grand unification scale, which cannot be directly tested. This comment also holds true when gauge non-singlet scalars ($SU(2)_{L}$ triplets in type-II seesaw [6, 7, 8]) or fermions ($SU(2)_{L}$ triplets in type-III seesaw [9]) are introduced to generate small neutrino masses.
@@
The Zee-Babu model [13, 14] extends the SM with two new scalars, $\eta^{+}(1,1,1)$ and $k^{++}(1,1,2)$, with $SU(3)_{C}\times SU(2)_{L}\times U(1)_{Y}$ quantum numbers as indicated. Their Yukawa couplings, $f_{ij}L_{i}L_{j}\eta^{+}+g_{ij}e^{c}_{i}e^{c}_{j}k^{--}$, with $f_{ij}=-f_{ji}$ and $g_{ij}=g_{ji}$ ($i,j$ are family indices here), as well as a cubic scalar coupling $\mu\,(\eta^{+}\eta^{+}k^{--})$ jointly break lepton number by two units, and would generate neutrino Majorana masses via two-loop diagrams involving the exchange of these scalar fields. An $SU(5)$ embedding of these fields can be realized by the addition of a $10_{H}$ and a $50_{H}$ to the scalar sector of minimal $SU(5)$ with antisymmetric $10_{H}$ and symmetric $50_{H}$ Yukawa coupling matrices, which is what our framework adopts. Note that the model has no right-handed neutrino $\nu_{R}$, which goes well with $SU(5)$ unification, which also has no $\nu_{R}$. The mass relations $m_{b}^{0}/m_{\tau}^{0}=m_{s}^{0}/m_{\mu}^{0}=m_{d}^{0}/m_{e}^{0}=1$ among the quarks and leptons predicted by the minimal $SU(5)$ model (the superscripts 0 indicate GUT scale values of the masses), while a good order of magnitude estimate, fail in detail when compared to experimental values by more than 100%. These relations can be corrected by introducing a vector-like pair of fermions in the $5+\overline{5}$ representation [24]. Gauge coupling unification is realized in this framework consistently, see Fig. 3, by choosing the mass scales of certain particles to be below the GUT scale, as listed in Table 2. A key outcome of our analysis is that, in addition to the original diagrams mediated by $\eta^{+}$ and $k^{++}$ scalars, neutrinos gain mass through diagrams mediated by their color-triplet and color-sextet counterparts [25, 26, 27], see Fig. 1. Once the experimental limits from proton decay are combined with gauge coupling unification constraints and perturbativity of Yukawa couplings, we find that both the color-neutral and the colored Zee-Babu states should have masses below $\mathcal{O}(10^{3})$ TeV, which can potentially be discovered in future collider experiments, thereby testing the proposed model.
Neutrino mass generation in the framework of $SU(5)$ unification has been studied in conjunction with gauge coupling unification by various authors. Ref. [28] has studied type-II seesaw embedding in $SU(5)$, while Ref. [29] has proposed type-III seesaw embedding. Models of radiative neutrino masses [15, 16] have also been embedded in $SU(5)$. The original Zee model [10], where the neutrino masses arise as one-loop radiative corrections, has been embedded in $SU(5)$ in Ref. [30, 31, 32]. For the embedding of other loop-induced neutrino mass models in $SU(5)$, see for example, Ref. [33, 34, 35, 36, 37, 38, 39, 40, 41] (see also Refs. [42, 43, 44, 45]). To the best of our knowledge the model presented here is the first attempt to embed the Zee-Babu model within $SU(5)$.
@@ 2 Embedding the Zee-Babu model in $SU(5)$
Once the GUT symmetry is broken, the submultiplets of the scalar fields will receive non-degenerate masses. In the context of a non-supersymmetric $SU(5)$ GUT, achieving gauge coupling unification while satisfying current proton decay bounds requires a split mass spectrum of the scalar fields. In the standard $SU(5)$ model, considering the adjoint Higgs, the color octet mass and the $SU(2)$ triplet mass are independent. See, for example, Eq. (8) of Ref. [47] (see also Ref. [48]). If a hierarchical mass spectrum is desired, one has to make a fine-tuning to get the weak triplet, for example, lighter than the color octet. Since the full Higgs potential has several parameters, it is possible to realize all the hierarchies that are needed for gauge coupling unification. Although there is no mass relations among the submultiplets of $5_{H},24_{H}$, and $10_{H}$, there exist three relations within the components of $50_{H}$ [49, 35], which read
@@
These mass relations will play crucial role in gauge coupling unification, which we discuss in Sec. 5. As for the doublet-triplet mass splitting, the situation in our case is similar to other non-supersymmetric $SU(5)$ models, which is achieved by fine-tuning.
@@ 3.1 Charged fermion mass
As will be shown later in Sec. 5, from the gauge coupling unification constraints, the vector-like lepton doublet should have a mass near the GUT scale. Therefore, we integrate it out, which yields the light $3\times 3$ charged lepton mass matrix and the heavy vector-like lepton mass as [50]
$M_{E}^{\mathrm{light}}=\left(\mathds{1}+\frac{1}{\xi_{E_{4}}^{2}}\xi_{E}\xi_{E}^{\dagger}\right)^{-1/2}v_{5}\hat{Y}_{5}^{\mathrm{diag}},$ (3.7)
@@
$M_{D}^{\mathrm{light}}=v_{5}\hat{Y}_{5}^{\mathrm{diag}}\left(\mathds{1}+\frac{1}{\xi_{D_{4}}^{2}}\xi_{D}\xi_{D}^{\dagger}\right)^{-1/2},$ (3.9)
@@ 3.2 A fit to the charged fermion masses
(at $\mu=M_{GUT}$) pull
$m_{d}/10^{-3}$ 1.14$\pm$0.11 1.14 0.021
$m_{s}/10^{-3}$ 21.58$\pm$1.14 21.58 0.064
$m_{b}$ 0.994$\pm$0.009 0.994 0.066
$m_{e}/10^{-3}$ 0.470692$\pm$0.000470 0.469544 0.146
$m_{\mu}/10^{-3}$ 99.3658$\pm$0.0993 99.2097 0.193
$m_{\tau}$ 1.68923$\pm$0.00168 1.68976 0.042
$\theta^{\rm{CKM}}_{12}$ $0.22739\pm 0.0006$ 0.2276 0.100
$\theta^{\rm{CKM}}_{23}/10^{-2}$ $4.858\pm 0.06$ 0.0486 0.058
$\theta^{\rm{CKM}}_{13}/10^{-3}$ $4.202\pm 0.13$ 0.00421 0.115
$\delta^{\rm{CKM}}$ $1.207\pm 0.054$ 1.203 0.067
Table 1: Inputs [51] and outputs for the numerical $\chi^{2}$ analysis. This benchmark fit with a total $\chi^{2}=0.1$ shows excellent agreement with data. As an example, for this fit, we additionally demand that the heavy down-type vectorike fermion mass is $m_{B}\sim 3$ TeV, which safely evades all present collider bounds.
@@
Specifically, in our numerical fit procedure, we minimize a $\chi^{2}$-function defined as
$\chi^{2}=\sum_{i}\mathrm{pull}_{i}^{2},\;\;\;\;\;\mathrm{pull}_{i}=\frac{T_{i}-O_{i}}{\sigma_{i}}.$ (3.11)
Here, $T_{i},O_{i},$ and $\sigma_{i}$ represent the theory prediction, experimentally observed central value, and the experimental error associated with the measurements, respectively, for the $i$-th observable. In the above equation, the sum is taken over all relevant observables to be fitted. As mentioned above, we fit the down-type quark and charged lepton masses. As will be discussed in the next sub-section, while checking for the flavor violating constraints, one must also need to determine the Cabibbo–Kobayashi–Maskawa (CKM) matrix. Therefore, in the $\chi^{2}$-function, we also include the quark mixing parameters. The input values in the CKM matrix are summarized in Table 1. Therefore, the sum in the $\chi^{2}$-function runs over three down-type quark masses, the mass of the bottom-like heavy quark $m_{B}$, three charged lepton masses, three mixing angles and the Dirac phase in the quark sector. As for the number of parameters, we have three real and positive couplings $y_{i}$, three ratios $\xi_{E_{i}}/\xi_{E_{4}}$ (which we take to be real), and four $\xi_{D_{a}}$ (for generality, three of which are taken to be complex). As discussed above, we also need to keep $\xi_{D_{4}}$ as a parameter to fit the mass of the heavy bottom-like quark. Moreover, the parametrization we use for the CKM-like matrix in the up-type quark sector, Eq. (3.20), has three mixing angles and three phases. Hence, we have 13 magnitudes and 6 phases to fit above-mentioned 11 observables (correspondingly, the number of degrees of freedom is 8). Note that we minimize the $\chi^{2}$-function to obtain a benchmark parameter set which satisfies all experimental constraints–we have no intention for obtaining the best fit. This provided benchmark solution has a total $\chi^{2}=0.1$ (hence, $\chi^{2}$ per number of degrees of freedom is about 0.01) and pulls of all observables are summarized in Table 1.
@@ 3.3 Flavor violation
where $V={V^{u}_{L}}^{\dagger}X_{L}$ is a $3\times 4$ matrix. Since there is no up-type VLQ, $V^{u}_{L}$ is a $3\times 3$ unitary matrix that diagonalizes the up-quark mass matrix. Therefore, the CKM matrix can be identified with the upper $3\times 3$ block of $V$, i.e., $V_{\mathrm{CKM}}=V^{\mathrm{upper}}_{3\times 3}$. During the fitting procedure, we also demand that the CKM matrix is correctly reproduced in a way that all non-unitarity constraints are satisfied. For this purpose, we parametrize $V^{u}_{L}$ as
$V^{u}_{L}=V^{u}_{\mathrm{CKM-like}}(\theta_{ij}^{u},\delta^{u})\mathrm{diag}(e^{i\alpha^{u}},e^{i\beta^{u}},1).$ (3.20)
@@
where we have defined $F^{d}=X_{L}^{\dagger}X_{L}$. Deviation of $F^{d}$ from the unit matrix gives rise to flavor changing neutral current (FCNC) processes, such as meson decay and meson-antimeson oscillations. The coupling responsible for Higgs-mediated FCNC takes the form [56]
@@
The fitted values of the CKM mixing parameters are also summarized in Table 1. Eqs. (3.12)-(3.14), together with Eq. (3.23), let us compute the relevant matrices as:
@@
$BR\big{(}\mathcal{B}_{q}\rightarrow\mu^{+}\mu^{-}\big{)}=\tau_{\mathcal{B}_{q}}\frac{G^{2}_{F}}{16\pi}\bigg{(}\frac{\alpha_{\text{em}}}{\pi s^{2}_{W}}\bigg{)}^{2}f^{2}_{\mathcal{B}_{q}}m_{\mathcal{B}_{q}}m^{2}_{\mu}\sqrt{1-\frac{4m^{2}_{\mu}}{m^{2}_{\mathcal{B}_{q}}}}|\eta^{2}_{Y}|\big{|}(V_{\mathrm{CKM}}^{*})_{tq}(V_{\mathrm{CKM}})_{tb}Y_{0}(x_{t})+\Delta^{\mathcal{B}_{q}}_{\mu\mu}\big{|}^{2}\,.$ (3.43)
In the above equation, $\tau_{\mathcal{B}_{q}}$ and $f_{\mathcal{B}_{q}}$ are the relevant lifetimes and decay constants, respectively. The factor $\eta_{Y}$ accounts for the QCD corrections [60] and $Y_{0}$ is one of the well-known Inami-Lim functions [61]. By demanding that the new contribution be smaller than the SM prediction at $95\%$ C.L, the corresponding constraints yield
@@
Finally, using the flavor violating couplings quoted above, one can straightforwardly compute the decay rates of the vectorlike quark [56], and we find
@@ 3.4 Neutrino mass generation
Although $\hat{Y}_{A}$ remains anti-symmetric, $\widetilde{Y}_{A}$ is no longer anti-symmetric. For the latter, this is not required by the SM gauge symmetry. Furthermore,
@@ 4 Perturbativity bounds on the Yukawa couplings
Before moving to the gauge coupling unification analysis, in this section, we study perturbativity bounds on the Yukawa couplings and find the upper bounds on the Yukawa couplings involving the Zee-Babu scalars. In particular, we choose the maximum values of Yukawa couplings at the low scale demanding that they remain perturbative (i.e., $y\leq\sqrt{4\pi}$) up to the GUT scale. For this analysis, here we assume that the color-neutral as well as the colored Zee-Babu states all have masses at the TeV scale. Then, the evolution of the gauge and the Yukawa couplings are governed by the one-loop renormalization group equations (RGEs) given by
@@ 5 Gauge coupling unification and proton decay bounds
In this section, we perform a dedicated study of the gauge coupling unification. The two-loop RGEs [69, 70, 71] for the gauge couplings can be written as
@@
Figure 4: Result from extensive Markov chain Monte Carlo (MCMC) analysis. The $1\sigma$ (dark colored band) and $2\sigma$ (light colored band) highest posterior density (HPD) intervals of the masses of all multiplets from gauge coupling unification constraints. Gray horizontal dashed line represents the approximate collider bound of $m_{i}\geq 1$ TeV on the states. Gray horizontal dotted line depicts the maximum allowed mass of $m_{i}\leq 10$ TeV on the $(\eta_{1},\chi_{1})$ states. Moreover, gray horizontal dashed-dotted line at $\mu=10^{12}$ GeV scale shows the proton decay bounds on the color-triplet states $\chi_{2}$ and $\phi_{2}$.
@@
Let us briefly discuss the fine-tuning needed, for example, for the case of the adjoint field. Following Eq. (8) of Ref. [47], one can write the masses of the weak-triplet, $\Phi_{2}$, and the color-octet, $\Phi_{3}$, as $m^{2}_{\Phi_{2}}=\left(-\mu+4\kappa v_{24}\right)v_{24}$ and $m^{2}_{\Phi_{3}}=\left(\mu+\kappa v_{24}\right)v_{24}$. Here, $\mu$ and $\kappa$ are relevant (properly re-scaled) coefficients of cubic and quartic terms in the scalar potential, see Ref. [47]. From the above analysis, it is clear that gauge coupling unification demands $m_{\Phi_{2}}\sim M_{\mathrm{GUT}}$, while $m_{\Phi_{3}}\sim$ TeV. Therefore, assuming natural value, in particular, $\mu\sim\mathcal{O}(M_{\mathrm{GUT}})$ and $\kappa\sim\mathcal{O}(1)$, we accept a fine-tuning of order $\mu+\kappa v_{24}\sim\mathcal{O}(\mathrm{TeV}^{2}/M_{\mathrm{GUT}})\sim 10^{-10}$ GeV, leading to $\mu\approx-\kappa v_{24}$. A similar analysis can be trivially extended for the rest of the multiplets.
@@
Finally, consistent gauge coupling unification points (from the above mentioned MCMC analysis) and proton decay bounds in the $g_{\mathrm{GUT}}$ vs $M_{\mathrm{GUT}}$ plane is depicted in Fig. 5. The corresponding proton decay bound is computed as follows. The decay rate for the dominant gauge boson mediated decay $p\rightarrow e^{+}\pi^{0}$ are given as [79, 80, 81]
@@
Here $f_{\pi}=130.2\,\mathrm{MeV}$ is the pion decay constant $D=0.804$ and $F=0.463$ are chiral Lagrangian parameters  [82, 83], $\alpha_{H}=-0.01257\,\mathrm{GeV}^{3}$ is the hadronic matrix element [84], $R_{L}=1.35$ is the long-distance renormalization factor of the four-fermion operators [85], $A_{SL}$ and $A_{SR}$ are the short distance factors with $A_{SL}\simeq A_{SR}\simeq 2$ [86], $M_{X}$ is the common mass of the $(X,Y)$ gauge bosons, and $\alpha_{\rm GUT}=g_{\rm GUT}^{2}/(4\pi)$ is the unified coupling at the GUT scale. Note that the value of $\alpha_{\rm GUT}\simeq 1/14$ is significantly larger than the corresponding value in typical non-supersymmetric unified theories where it is $\sim 1/40$. This increased $\alpha_{\rm GUT}$ results in the shortening of proton lifetime by about an order of magnitude. For the gauge coupling unification study, we have identified the unification scale $M_{\mathrm{GUT}}$ with the masses of the gauge bosons, i.e., $M_{\mathrm{GUT}}=M_{X,Y}$. Using Eq. (5.62), in Fig. 5, the current (future) proton decay bound of $\tau_{p}(p\to e^{+}\pi^{0})>2.4\times 10^{34}$ ($7.8\times 10^{34}$) yrs from Super-Kamiokande [87] (Hyper-Kamiokande [23]) is shown with gray shaded region (red dashed line). While, a large part of the parameter space is ruled out by the present experimental limits, intriguingly, 10 yrs operation of Hyper-Kamiokande will likely probe the entire available parameter space of the proposed model. Although there are more gauge mediated proton decay modes, the channel $p\to e^{+}\pi^{0}$ is the most dangerous and provides the most stringent limits. Additionally, color-triplet scalars also mediate proton decay. As discussed above, we ensure that all relevant color-triplets have masses $\gtrsim 10^{12}$ GeV (see, for example, Refs. [72, 39]) so that corresponding experimental limits are satisfied.